OpenStudy (anonymous):

can someone help me? use a matrix to find a solution to the system of equations.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

start by dividing eq.1 by 8 ((1/8)R1) -x - y = -2 6x - 9y = -108 6R1 + R2: -x - y = -2 -15y = -120 divide both sides of eq.2 by -15: y = 8 use back substitution: -x -8 = -2 -x = 6 x = -6 (-6,8) so the answer is a. :P

OpenStudy (anonymous):

thank you so much!!! omg

OpenStudy (anonymous):

No Problem :)